OpenStudy (anonymous):

reduction of order yy'-y''=0

OpenStudy (anonymous):

Never had a course in differential equations. Will the following help?\[y(x)=\sqrt{2} \sqrt{c_1} \tan \left(\frac{\sqrt{c_1} \left(c_2+x\right)}{\sqrt{2}}\right) \]